The cost of renting a car for a day is given by the linear function c(x) = 1/8 + 20 where c(x) is in dollars and x is the number of miles driven.
1. What is the cost of driving the care 180 miles?
2. How many miles could you drive if you had a budget of $100.
3. Explain the meaning of the 1/8 in the equation and what it represents in this context.
4. Explain the meaning of the 20 in the equation and what it represents in this context.
5. Graph c on the axes and label at least 3 points on a graph.

Answer by mukhopadhyay(490)

I think the function will be c(x) = (1/8)x + 20
1. c(180) = 180/8 + 20 = 45/2 + 20 = $42.50
2. 100 = x/8 + 20 => x=8(80) => x=640
3. It represents the variable cost per mile; varies by number of miles driven.
4. It represents a flat charge that the renter must pay, even if the car is not driven at all.
5. This is a linear function. Pick any two x's and find c(x) for those x's. Connect the dots as a staright line.